Show that the tangent from the point with vector -2i-3j to the ellipse 4x^2 +9y^2=36 are perpendicular?
1
Expert's answer
2020-03-10T12:53:34-0400

"ellipse :4x^2 +9y^2=36"

Converting this into standard form

it becomes : "\dfrac{x^2}{9}+\dfrac{y^2}{4}=1"

"a^2=9 \ \&\ b^2=4"


Given vector : "-2i-3j"

Converting this into point it becomes (-2,-3)

Let the tangent to the ellipse from point "y=mx+c"

then it must satisfy the condition "c^2=a^2m^2+b^2"

substituting the values in both the equation we get :

"-3=-2m+c" ……..(1) & "c^2=9m^2+4" ……..(2)

Putting the value of c in the eq(2) from the equation 1

"(2m-3)^2=9m^2+4\\4m^2+9-12m=9m^2+4\\5m^2+12m-5=0\\"

From here we will get two values of m(slope)

If the tangents are perpendicular then the product of slopes must be -1

product of roots of the above quadratic = -5/5=-1